On June 1, 2020 Night sold land to Blue Company in exchange for a $700,000 non-interest bearing note due on June 1, 2030. The prevailing rate of interest for a note of this type was not available. The cost of land to Night was $250,000. Night would have accepted $390,876 in cash for the land. Answer the following questions with whole numbers.

1) Calculate the gain in sale that Night will record from the sale of the land.

2) Indicate the interest rate implicit in the note receivable.

3) Calculate the interest revenue in the December 31, 2020 income statement

4)Indicate the carrying value of the note receivable reported on the December 31, 2020 balance

5) indicate the total interest revenue that Nigh will record over the ten year life of the note

Please show all the work and numbers used so I can compare against the answer.
Answer #1
Gain On sale that Night will record from the sale of the land
Acceptance of cash for the land        3,90,876
Less: Cost of the land        2,50,000
Gain on sale of land        1,40,876
interest rate implicit in the note receivable.
PV of $ 700000 at x% for 10 years = 390876
The rate will be 6%
700000 * (PVIF 6%,10) = 700000 * 0.558395 = 390876
interest revenue in the December 31, 2020 income statement
Interest revenue = 390876 * 6% * 7/12 = 13681
The carrying value of the note receivable reported on the December 31, 2020 balance will be 390876
total interest revenue that Nigh will record over the ten year life of the note
Face value 700000
Less: Present Value 390876
total interest revenue that Nigh will record over the ten year life of the note
309124